Two joggers start from opposite ends of an 8 mile course running towards each other. One jogger is running at a rate of 4 mph, and the other is running at a rate of 6 mph where will they cross paths? Write a system of inequalities to represent the situation.


Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Two joggers start from opposite ends of an 8 mile course running towards each other. One jogger is running at a rate of 4 mph, and the other is running at a rate of 6 mph where will they cross paths?
-------
Jogger #1 DATA:
rate = 4 mph ; time = t hrs; distance = 4t miles
-----
Jogger #2 DATA:
rate = 6 mph ; time = t hrs ; distance = 6t miles
-------
Equation:
4t + 6t = 8 miles
10t = 6 miles
time = 0.6 hrs = 0.6*60 min = 36 minutes
====
Where will they cross paths?
From Jogger #1 start point:: (36/60)4 = 36/15 = 2.4 miles
From Jogger #2 start point:: 8-2.4 = 5.6 miles
-----------------
Cheers,
Stan H.
----------------

Answer by MathTherapy(10556)


Two joggers start from opposite ends of an 8 mile course running towards each other. One jogger is running at a rate of 4 mph, and the other is running at a rate of 6 mph where will they cross paths? Write a system of inequalities to represent the situation.
Slower jogger (4 mph) meets 
Faster jogger (6 mph) meets